### 17 Feb 2026 v3

# Two-dimensional RMSD projections for reaction path visualization and validation

Abstract
Transition state or minimum energy path finding methods constitute a routine component of the computational chemistry toolkit. Standard analysis involves trajectories conventionally plotted in terms of the relative energy to the initial state against a cumulative displacement variable, or the image number. These dimensional reductions obscure structural rearrangements in high dimensions and are often history dependent. This precludes the ability to compare optimization histories of different methods beyond the number of calculations, time taken, and final saddle geometry. We present a method mapping trajectories onto a two-dimensional projection defined by a permutation corrected root mean square deviation from the reactant and product configurations. Energy is represented as an interpolated color-mapped surface constructed from all optimization steps using a gradient aware derivative Gaussian Process. This representation highlights optimization trajectories, identifies endpoint basins, and diagnoses convergence concerns invisible in one-dimensional profiles. We demonstrate the framework on a cycloaddition reaction, showing that a machine-learned potential saddle and density functional theory reference lie on comparable energy contours despite geometric displacements, along with the ratification of the visualization for more complex reactions, a grignard rearrangement, and a bicyclobutadiene rearrangement.
Keywords Nudged Elastic Band, Saddle Search Methods, Visualization, Transition States
The first order saddle point, is often the first approximation to understand the reaction kinetics of any system of interest. Starting from two known configurations, a series of configurations connected through fictitious springs, that is, the nudged elastic band method [18] is among the most common. The NEB may be coupled with the climbing image [16] and spring variations [1] with machine learned local accelerations [14, 20, 11] and the string family of methods [28, 34, 24]. Together, these form the current Pareto optimal forms to find a saddle.
Beyond difficulties of convergence and the intricacies of the initial points chosen, the resulting path histories are best studied through the “eye-ball norm”, wherein manual inspection of the final path, along with a normal mode analysis for the saddle point estimate are considered sufficient. In turn, this means the most common visual plots, are one dimensional “profile” plots. These may show only the final optimized path, against the image number 111e.g. as found in ASE [21], which obscures any distance measure between the images. Alternatively, a well-known improvement to this involves the cubic Hermite interpolation [16] involving the forces relative to the “reaction coordinate,” defined by the piece-wise sum of the Euclidiean distances between intermediate images. The reaction coordinate $s_{i}=\sum_{j=1}^{i}\|\mathbf{R}_{j}-\mathbf{R}_{j-1}\|_{2}$ depends entirely on the specific path geometry and optimization history. This scalar measure lacks a unique definition and collapses global geometric information onto a single, arbitrary axis. This dimensional reduction precludes rigorous comparison between trajectories generated by differing algorithms, such as NEB versus Frozen String methods, or even identical algorithms with varying parameters. Furthermore, the one-dimensional projection frequently masks the distinction between numerical instability and physical relaxation into alternative basins, creating ambiguity in the validation of stationary points which costs researchers often days of calculations to clarify.
1 Methodology
Figure 1: Overview of the 2D RMSD projection pipeline. NEB output in $\mathbb{R}^{3N}$ is projected to intrinsic $(r,p)$ coordinates via IRA-based permutation-invariant RMSD. Tangential forces are projected onto the 2D tangent to construct synthetic gradients. Energies and gradients jointly feed a gradient-enhanced GP with the IMQ kernel to produce the interpolated energy surface $E(r,p)$.
The complete methodology is summarized in Figure 1. We describe the one-dimensional profiles first. Since standard NEB implementations provide discrete images $\mathbf{X}_{i}$ and energies $E_{i}$. To reconstruct a physically consistent energy profile $E(s)$ along the path, we utilize the available force information. We define the discrete reaction coordinate $s_{i}$ as the cumulative Euclidean distance:
$s_{i}=\sum_{j=1}^{i}\|\mathbf{X}_{j}-\mathbf{X}_{j-1}\|$ (1)
Instead of simple linear or cubic interpolation, we employ a Piecewise Cubic Hermite Interpolating Polynomial (PCHIP). This method utilizes the tangent forces $F_{\parallel,i}$ to constrain the derivative of the energy surface:
$\frac{dE}{ds}\bigg|_{s_{i}}=-F_{\parallel,i}=-(\mathbf{F}_{i}\cdot\hat{\tau}_{i})$ (2)
where $\hat{\tau}_{i}$ denotes the unit tangent vector along the path.
1.1 Intrinsic Projection Coordinates
To map the high-dimensional optimization trajectory onto a consistent 2D subspace, we define the projection coordinates $(u,v)$ as the distances from the reactant ($\mathbf{R}$) and product ($\mathbf{P}$) reference configurations. To ensure the metric remains robust across automated workflows where atom indexing may vary, we employ a permutation-invariant Root Mean Square Deviation (RMSD) with optimal permutation $\mathbf{P}$ and rotation $\mathbf{R}$ determined via the Iterative Rotations and Assignments (IRA) [15] algorithm. This procedure guarantees unique, invariant coordinates regardless of the initial atom indexing or frame orientation [11]. The discrete optimization steps provide a sparse sampling of this geometric subspace.
For a system with $N$ atoms having positions $\mathbf{X}\in\mathbb{R}^{3\times N}$, we define the distance metric $d(\mathbf{X},\mathbf{X}_{\text{ref}})$ as:
$d(\mathbf{X},\mathbf{X}_{\text{ref}})=\min_{\mathbf{Q},\mathbf{\Pi}}\sqrt{\frac{1}{N}\left\|\mathbf{X}-\mathbf{Q}\mathbf{X}_{\text{ref}}\mathbf{\Pi}\right\|_{F}^{2}}$ (3)
Here, $\|\cdot\|_{F}$ denotes the Frobenius norm. $\mathbf{Q}\in SO(3)$ represents the optimal rotation matrix, and $\mathbf{\Pi}$ represents the optimal permutation matrix. We solve for $\mathbf{Q}$ and $\mathbf{\Pi}$ simultaneously. This ensures that the resulting coordinates $(u,v)=(d(\mathbf{X},\mathbf{R}),d(\mathbf{X},\mathbf{P}))$ remain invariant to rigid body rotation, translation, and arbitrary atom index labeling.
[…]
For systems with well-established collective variables (Ramachandran angles for peptides, donor-acceptor distances for proton transfer), such coordinates naturally provide superior physical insight [22]. However, the vast majority of transition state searches lack such convenient coordinates. Surface reactions, where an adatom hops into a subsurface site on a metallic slab, or catalytic processes involving concerted rearrangements of adsorbates, have no obvious low-dimensional collective variable. Constructing a two-dimensional free energy surface for such systems requires extensive sampling and a priori knowledge of the relevant degrees of freedom. In contrast, the RMSD-based projection requires only the Cartesian coordinates already produced by the calculation. Unlike free energy surface methods that mandate pre-specification of reaction coordinates, our approach visualizes arbitrary path based methods post hoc, requiring only the endpoints. This eliminates the circular dependency where one must already understand the mechanism to choose coordinates capable of revealing it.
Our method targets the complementary regime to global PES exploration techniques: we provide geometric validation for individual optimization trajectories ($\approx 10^{3}$ samples per NEB calculation) rather than exhaustive basin catalogs requiring $\approx 10^{6}-10^{9}$ samples. Most single NEB calculations involve only $\approx 10^{3}$ geometries concentrated along a one-dimensional path connecting a single reactant-product pair. This sampling is insufficient and inappropriate for techniques designed to map complete basin connectivity or construct disconnectivity graphs [33]. Our framework instead answers a distinct question: “Did this specific optimization find a physically reasonable barrier topology?” This diagnostic capability proves critical for MLIP validation, where the goal is verifying that a machine-learned potential reproduces the correct transition pathway, not exhaustively cataloging all possible pathways.
1.2 Energy Landscape Projection
Figure 2: Dimensional reduction from the full 3N13N{-}1 configuration space to 1D (standard reaction coordinate $s$, left) versus 2D RMSD projection (this work, right). Both maps are many-to-one: distinct Cartesian configurations can share the same projected coordinates. The 1D projection loses all geometric context. The 2D projection retains endpoint distances, enabling cross-method comparison and landscape visualization, at the cost of 3N33N{-}3 unconstrained degrees of freedom.
The discrete optimization steps provide a sparse, unstructured sampling of the 2D $(u,v)$ domain (Figure 2). To visualize slices of the underlying potential energy landscape $E(u,v)$, in a meaningful way to compare data from disparate calculations, we construct a continuous projection approximating the potential energy surface. Both the 1D and 2D projections are lossy: multiple distinct configurations in $\mathbb{R}^{3N}$ can map to the same projected coordinate. The 2D projection retains strictly more information than the 1D reaction coordinate, but it is not and cannot be the full PES.
While forces $F=-\nabla_{R}(E)$ are available from electronic structure calculations in 3N-dimensional Cartesian space, direct projection via the chain rule would require computing $\partial d_{\text{RMSD}}/\partial R$, a quantity complicated by the minimization over permutations and rotations implicit in the RMSD definition. Instead, we construct synthetic gradients in the 2D projection space by combining the path tangent vector with the available parallel force component.
First, we stabilize the tangent calculation via Savitzky-Golay smoothing of the RMSD coordinates themselves:
$\tilde{r}_{i},\tilde{p}_{i}=\text{SavGol}(r_{i},p_{i};\text{window}=5,\text{poly}=2)$ (4)
From the smoothed coordinates, we compute the path tangent:
$\tau_{r}=\frac{d\tilde{r}}{ds},\quad\tau_{p}=\frac{d\tilde{p}}{ds}$ (5)
where the tangent is normalized so $\tau||=1$. The projected gradient components are then:
$\nabla_{r}E=-F_{\parallel}\cdot\tau_{r},\quad\nabla_{p}E=-F_{\parallel}\cdot\tau_{p}$ (6)
where $F_{\parallel}$ is the force component parallel to the path (column 3 in an eOn .dat file). This construction ensures gradient information respects the local path geometry while remaining tractable to compute from standard NEB output.
Only the tangential force component $F_{\parallel}$ enters this construction. The full Cartesian force $\mathbf{F}_{i}\in\mathbb{R}^{3N}$ contains orthogonal components encoding local curvature of the true PES. However, projecting these into the $(r,p)$ subspace requires the Jacobian $\partial d_{\text{RMSD}}/\partial\mathbf{R}$, which is not analytically available: the RMSD involves a joint minimization over rotations $\mathbf{Q}\in SO(3)$ and permutations $\mathbf{\Pi}$, making the map from $\mathbb{R}^{3N}$ to $\mathbb{R}^{2}$ piecewise smooth at best. The tangential projection is therefore the consistent choice without additional approximations. It still carries non-trivial 2D information because the tangent direction $(\tau_{r},\tau_{p})$ varies along the path, distributing force information across both coordinates in a way not accessible from scalar reaction coordinate plots.
We stress that the interpolated surface is a projected slice of the full \$3N\$-dimensional PES: many distinct Cartesian configurations map to the same $(r,p)$ point. The interpolation represents the conditional expectation of the energy given the RMSD coordinates, marginalizing over the unconstrained 3N23N-2 degrees of freedom. Features of the projected surface, such as apparent alternative pathways in the color map, do not necessarily correspond to physical pathways in the full configuration space. Similarly, the property that MEP images are minima in the orthogonal hyperplane holds in the full-dimensional PES but has no analogue in a two-dimensional lossy projection. The interpolation is reliable near the sampled data and degrades with distance, as with any regression model.
We denote the 2D projection coordinates compactly as $\mathbf{x}_{i}=(r_{i},p_{i})=(d_{\text{RMSD}}(S_{i},R),\,d_{\text{RMSD}}(S_{i},P))$. Together with the projected gradients $(\nabla_{r}E,\nabla_{p}E)$ from above, these form the inputs to a Gaussian process regression with derivative observations [13, 23, 9]. We construct the energy surface using the Inverse Multiquadric (IMQ) kernel:
$k(\mathbf{x},\mathbf{x}^{\prime})=\left(c^{2}+r^{2}\right)^{-1/2},\quad r^{2}=\|\mathbf{x}-\mathbf{x}^{\prime}\|^{2}$ (7)
where $c$ is a scale parameter. The function $g(t)=(c^{2}+t)^{-1/2}$ is completely monotone on $0,\infty)$: $(-1)^{n}g^{(n)}(t)\geq 0$ for all $n\geq 0$. By Schoenberg’s theorem [30], any kernel of the form $k=g(\|\mathbf{x}-\mathbf{x}^{\prime}\|^{2})$ with $g$ completely monotone is strictly positive definite on $\mathbb{R}^{d}$ for all $d$ [6]. This guarantees non-singular interpolation matrices regardless of point configuration.
The gradient-enhanced GP requires derivative kernel blocks. For the IMQ these are:
$\frac{\partial k}{\partial x_{i}}$ $=-(x_{i}-x_{i}^{\prime})\,(c^{2}+r^{2})^{-3/2}$ (8)
$\frac{\partial^{2}k}{\partial x_{i}\partial x_{j}^{\prime}}$ $=\delta_{ij}(c^{2}+r^{2})^{-3/2}-3(x_{i}-x_{i}^{\prime})(x_{j}-x_{j}^{\prime})(c^{2}+r^{2})^{-5/2}$ (9)
which assemble into the augmented kernel matrix [19] over the observation vector $\mathbf{y}_{\text{full}}=[E_{1},\nabla_{r}E_{1},\nabla_{p}E_{1},E_{2},\ldots]$. In practice, the derivative covariances are computed via automatic differentiation using the JAX library, so extending to other kernels requires only changing the base kernel function.
As $r\to\infty$, the IMQ kernel decays as $r^{-1}$, its first derivatives as $r^{-2}$, and second derivatives as $r^{-3}$. This polynomial decay contrasts with the exponential decay of the Matérn and squared exponential families. Heavy tails allow the kernel to capture long-range basin structure from sparse path data, and the augmented kernel matrix remains better conditioned than exponentially decaying counterparts. Including derivative observations effectively triples the information content per sampled geometry without additional energy evaluations.
The hyperparameter optimization follows the subset optimization [12] concept from the OT-GPD, here, the points from the final path are used to calculate the length and noise scales, which are subsequently applied while fitting the entire surface. This subsampled approach reduces computational cost from $O(N^{3})$ to $O(n^{3})$ where $n\ll N$ (typically $n\approx 20$ vs $N\approx 500-2000$), while simultaneously improving robustness by isolating hyperparameter learning from transient optimization dynamics.
The resulting interpolated surface enables direct overlay of reference structures (e.g., DFT-optimized saddle points) onto MLIP-generated landscapes. Visual assessment of whether a potential captures qualitatively correct barrier topology becomes possible even when geometric displacements occur, a capability impossible with 1D energy profiles where the reaction coordinate axis itself depends on path geometry. This addresses a critical gap in current MLIP validation workflows.
2 Results
[…]
We validate the framework, first contrasting against 1D profiles for a the 1,3-dipolar cycloaddition of ethylene and N2O forming 4,5-dihydro-1,2,3-oxadiazole, a well-studied benchmark reaction for NEB method development [20, 1, 3, 24, 29, 5]. Figure 1 compares conventional one-dimensional representations with the two-dimensional RMSD projection for energy-weighted NEB optimization in eOn [10] 222from https://eondocs.org using the PET-OMAT machine-learned potential [25, 4] using Metatomic [3]. The model has been trained on PBE reference data from the Open Materials dataset [2], with full construction details in the associated publication. For the cycloaddition, we contrast this with a saddle optimized from a different NEB calculation using the B3LYP functional in ORCA [1] projected onto the interpolated landscape.
The conventional profiles (Fig.1, bottom) show optimization from an initial barrier of approximately 1.1 eV to a final value near 0.4 eV over 120 steps, with the product lying approximately 0.8 eV below the reactant. The final path appears smooth and well-converged. These representations provide no information about sampling quality or the broader landscape topology.
The two-dimensional RMSD projection (Fig.1, top) provides some more insight. First, sampled structures (black dots) cluster tightly along the converged path, confirming robust single-pathway convergence. Second, the interpolated energy surface displays clear landscape topology: a barrier region (yellow/orange, approximately 0.4eV) separating the reactant basin (upper-left) from the deeper product basin (blue, lower-right). The projection reveals that the ORCA saddle configuration overlaps with the estimate from the MLIP.
(a) Grignard Rearrangement
(b) Bicyclobutadiene Isomerization
Figure 4: Two-dimensional RMSD projections of complex rearrangement pathways. (a) The Grignard rearrangement of (Z)-2-phenyl-2-((trimethylsilyl)oxy)hex-4-enenitrile exhibits a distinct ”turning” mechanism. The projection reveals significant optimization instability (scattered black dots) near the reactant basin that 1D profiles would obscure. The star indicates the reference saddle from Birkholz et al. (b) The conrotatory ring opening of bicyclobutadiene shows a sharp geometric ”kink” post-transition state. The projection confirms the machine-learned potential path (circles) faithfully traverses the saddle region identified by the Free String Method (star).
To demonstrate robustness on non-trivial topologies, we analyze the Grignard rearrangement of (Z)-2-phenyl-2-((trimethylsilyl)oxy)hex-4-enenitrile [5] and the electrocyclic ring opening of bicyclobutadiene (Figure 2) [24]. Unlike the linear character of the cycloaddition, these reactions proceed through significant geometric turns. The Grignard pathway (Figure 2a) exhibits a curved trajectory where the path tangent rotates nearly 90 degrees in the $(RMSD_{R},RMSD_{P})$ plane. The 2D projection clearly captures the optimization noise (scattered black points) in the solvent-mediated reactant region, a diagnostic feature lost in scalar reaction coordinate plots, while demonstrating without explicitly needing to study the structures, the fact that they lie on the same energy contour. Similarly, the bicyclobutadiene isomerization (Figure 2b) displays a sharp geometric “elbow” immediately following the transition state. The interpolated surface correctly places the reference Free String Method (mlFSM) saddle [24] within the barrier region of the MLIP. This confirms that the NEB trajectory found the physically relevant saddle despite the complex topology, a validation that remains ambiguous when observing only the image index or scalar reaction coordinate 333or in practice, would involve normal mode analysis, barrier heights, RMSD, and visual inspection.
To be explicit about what the 2D projection provides beyond standard scalar comparisons: computing energy differences and RMSD values between two saddle points tells one that they differ, but not how they differ relative to the reaction landscape. A 1D energy profile, by construction, collapses all geometric information onto the cumulative displacement axis, making it impossible to distinguish whether two saddle estimates from different methods lie on the same energy contour, sit in different basins, or whether the optimization explored alternative pathways before converging. The 2D projection preserves these geometric relationships. One can immediately see, for example, that the ORCA and MLIP saddles in Figure 1 occupy the same barrier region despite different path histories, or that the optimizer noise in the Grignard case (Figure 2a) is confined to a specific geometric region rather than distributed along the path. These are qualitative assessments, analogous to plotting a trajectory on a map rather than reporting only the total distance traveled.
Furthermore, this framework provides a unified coordinate system for validating potential energy surfaces against higher-level theory. If these disagree, the relative differences between the saddles are apparent. The projections reveal whether resulting structures sit within comparable energy contours. Such an analysis, checking if the energy surface captures the qualitative barrier topology even if the precise saddle geometry differs due to functional sensitivity, is impossible with conventional one-dimensional profiles.
For complete reproduction, the materials archive includes both exact commands, and outputs at the time of submission.
Limitations
The $(r,p)$ projection is a lossy map from $\mathbb{R}^{3N}$ to $\mathbb{R}^{2}$. Multiple distinct Cartesian configurations can share the same RMSD coordinates, so the interpolated energy surface does not reconstruct the true PES but rather a projected slice of it. Apparent features far from the sampled data, for instance, alternative low-energy pathways visible in the color map, may not correspond to physical pathways in the full configuration space. The interpolation should be read as a qualitative guide near the data, not as a quantitative surrogate for the potential.
Because only the tangential force component enters the gradient projection, the method does not recover curvature information orthogonal to the path. In the full \$3N\$-dimensional PES, MEP images are by definition minima in the hyperplane orthogonal to the path. This property cannot be enforced or verified in the 2D projection without the Jacobian $\partial d_{\text{RMSD}}/\partial\mathbf{R}$, which is unavailable in closed form due to the rotation-permutation minimization. Any extension to orthogonal force projection would require differentiable structure matching, which is an open problem.
The choice of IMQ kernel, while theoretically motivated by its positive definiteness and polynomial tail properties, is one of several valid options. The software provides squared exponential, Matern 5/2, and thin plate spline alternatives. Different kernels may be better suited to different reaction topologies, and no single kernel is universally optimal.
Finally, the method is designed for post-hoc visualization of existing calculations. It does not accelerate the NEB computation itself, nor does it replace quantitative analysis (normal mode analysis, barrier heights, vibrational frequencies) required for kinetic rate calculations.
Conclusion
We presented a coordinate-free visualization method for analyzing reaction path optimization trajectories. By projecting high-dimensional pathways onto a surface defined by permutation-invariant RMSD coordinates, the method reveals geometric and energetic features obscured by standard one-dimensional profiles.
In current practice, comparing NEB results across methods or potentials requires either pointwise structural inspection (impractical beyond a handful of images) or scalar summaries (barrier height, RMSD to reference) that discard geometric context. The 2D projection provides a middle ground: a single plot that preserves the geometric relationship between all sampled structures, the converged path, and any reference points, while remaining interpretable at a glance. The benchmarks demonstrate that this representation distinguishes numerical artifacts from physical topology and reveals whether different potentials produce qualitatively similar barrier regions, even when the precise saddle geometries differ.
While we focus on double-ended methods here, the $(RMSD_{R},RMSD_{P})$ projection operates independently of the path generation algorithm. One can project histories from single-ended saddle search methods, molecular dynamics, or meta-dynamics onto these intrinsic axes post hoc to diagnose path quality and basin hysteresis. The relative energy surfaces from multiple machine learning models could also be compared within such a framework.
Conflict of Interest
We declare no conflicts of interest.